# I-Scene: 3D Instance Models are Implicit Generalizable Spatial Learners

### v1

@@ 1 Introduction
Generalization remains the central obstacle to interactive 3D scene generation. An effective system must produce editable, affordance‑aware, and spatially coherent object arrangements, capabilities critical for virtual content creation, simulation, and embodied AI. Recent end‑to‑end approaches extend powerful image‑to‑3D instance priors [xiang2025structured, tochilkin2024triposr] to multiple objects and relations in a single pass, yet their spatial understanding is typically learned from curated scene datasets whose limited coverage does not scale and ultimately constrains generalization to new layouts.
Recent methods learn layouts directly from annotated scenes [yang2024physcene, tang2024diffuscene, meng2025scenegen, huang2025midi]; for example, SceneGen [meng2025scenegen] explicitly models poses on 3D‑FRONT [fu20213d], and MIDI [huang2025midi] learns inter‑object relations from those annotations. However, available interactive scene datasets are limited in scale, diversity, and spatial variety. For instance, the widely used interactive 3D scene dataset 3D-FRONT contains only $\sim$20K indoor bedroom and living-room scenes and under-represents small and supporting objects. Consequently, scene layout‑supervised models often overfit to dataset‑specific biases and fail to generalize to broader scene distributions—for example, small objects placed on or behind large furniture or outdoor settings.
@@ 2 Related Work
Single geometry 3D scene generation. One line of work represents entire scenes as a single global structure guided by text or visual prior. For example, Wonderland [liang2025wonderland] predicts 3D Gaussian Splatting in a feed-forward manner from a single image; WonderWorld [yu2025wonderworld] generates connected scenes from a single view using layered Gaussian surfels. Similarly, WorldExporer [schneider_hoellein_2025_worldexplorer]
@@
Compositional 3D interactive scene generation. Recent pipelines decompose scenes into perception and assembly: instance detection/segmentation and depth provide amodal cues, then objects are retrieved or generated, finally are assembled into a scene. Gen3DSR [ardelean2024gen3dsr] follows a modular divide-and-conquer strategy; Deep Prior Assembly [zhou2024zero] integrates frozen priors (e.g., SAM [kirillov2023segment], diffusion, Shap-E [jun2023shap]) with layout fitting; REPARO [han2024reparo] performs compositional asset generation with differentiable 3D layout alignment; CAST [yao2025cast] applies component-aware analysis with SDF-based physics correction. On the text-driven side, graph/LLM planners (e.g., LayoutGPT [feng2023layoutgpt], Holodeck [yang2024holodeck], GALA3D [zhou2024gala3d]) synthesize scene graphs or relational constraints that downstream stages realize with assets and solvers. Scenethesis [ling2025scenethesis] and SceneWeaver [yang2025sceneweaver] combines LLM planning, vision guidance, and physics-aware optimization with a plausibility judge. These approaches provide high fidelity and open-ended semantics, but are sensitive to early perception/planning errors, potential retrieval/solver mismatch, and incur per-scene optimization, limiting throughput and scalability.
Learning-based 3D interactive scene generation. Most recent approaches learn spatial layout distributions from curated 3D scene datasets [fu20213d], differing mainly in how they instantiate object assets. i) Layout-first scene synthesis. model a room as an unordered set of object tokens with attributes such as category, size, pose, which are later instantiated as geometry via retrieval or generation. For instance, ATISS [paschalidou2021atiss], SceneFormer [wang2021sceneformer]) DiffuScene [tang2024diffuscene], MiDiffusion [hu2024mixed], DeBaRA [maillard2024debara], and PhyScene [yang2024physcene] condition on room type or floor plan, enabling controllable scene completion and synthesis; ii) End-to-end multi-instance method. With the advancement of pre-trained image-to-3D object model, recent approaches jointly model multiple objects and their spatial relations, reducing cascade errors and avoiding retrieval/solver loops at test time. MIDI-3D  [huang2025midi] and SceneGen [meng2025scenegen] model multiple assets and their implicit inter-object relations or instance explicit poses in one feed-forward pass. PartCrafter [lin2025partcrafter] extends compositional latent diffusion transformers to jointly denoise parts/objects into explicit triangle meshes. Both paradigms achieve strong control and coherence but remain constrained by the biases and limited diversity of existing scene datasets, often struggling with rare layouts or complex arrangements such as small objects supported by or occluded behind larger ones.
@@ 3.2 Overview
Figure 2 presents the overview of our generalizable spatial learner(I-Scene). I-Scene reprograms an image-to-3D instance foundation model to a spatial leaner. We use TRELLIS [xiang2025structured] backbone in the experiment. I-Scene only modifies the sparse structure transformer. Other stages are kept the same. I-Scene has two branches: spatial guidance branch and instance generation branch. The two branches share the same weight and are trained jointly.
@@ 3.4 View-Centric(VC) Space
Existing relevant works [huang2025midi, meng2025scenegen] follow image-to-3D base model and use canonical space for scene representation. We observe using canonical space limits the learning of $F$ and hurts the model generalization ability. As illustrated in Figure 3 (a), given two camera ($C1$ and $C2$) and the scene composed of a blue square and a red circle in the space, the camera views of the scene are different as they are in different location and orientation, but the instance in the scene stay at the same position in the canonical space. So this formulation makes the $F$ ignores the object spatial position in the view image $I$ and only focuses on the local shape of each instance in the image. When training on some indoor dataset like 3D-FRONT, it is not significantly hurting $F$ as there are limited object in the scene and each instance usually has very different shapes. But if the scene has identical objects, e.g. several identical chairs having similar poses, $F$ often put duplicated chairs at the same position. We argue that the object spatial layout in the RGB image is also important and provides strong hint on its 3D layout pose. So we propose view-centric space where the space axis is based on the camera pose. As shown in Figure 3 (a), given the same two camera setting, the spatial layout of the two objects follows the change of the camera pose coherently. When training I-Scene in VC space, $F$ achieves much better generalization results as shown in our ablation study(Table 3).
@@ 3.5 Non-Semantic Synthetic 3D Scene
Collision-free Random Layout. We generate our synthetic scenes via sampling high‑quality 3D instances from a diverse 3D asset dataset(e.g., Objaverse [deitke2023objaverse]), randomly places them with a collision‑free mechanism to reduce severe occlusions, and enforces a common spatial relation such as right, left, front, back, and on the top of. The scene layout is purely non‑semantic meaningful composition of instance; only geometric and basic physical plausibility constraints are applied. By training on this dataset, I-Scene learns general spatial reasoning while remaining agnostic to category semantics. We provide our random layout generation details in appendix.
@@ 4.1 Experiment Setting
Baseline. We compare the representative interactive 3D scene generation methods including MIDI [huang2025midi], SceneGen [meng2025scenegen], Partcrafter [lin2025partcrafter], and Gen3DSR [ardelean2024gen3dsr].
For all baselines, we use the same scene RGB image and instance masks as inputs, except PartCrafter, which does not support mask control; for PartCrafter we provide the scene RGB plus the target instance count. Because PartCrafter lacks texture rendering, we evaluate it only on scene-level geometry and layout quality. For visual quality comparisons, we texture MIDI’s generated scenes using MV-Adapter [huang2025mv].
@@
Evaluation Dataset. We evaluate all methods on synthetic data and real-world/stylized scenes. i) Synthetic (in-domain). Following MIDI, we adopt the 3D-FRONT split and use the same test scenes, filtering out renders with fully occluded instance masks; this yields $\sim$860 test scenes. ii) Synthetic (out-of-domain). To assess generalization to novel layouts and object sets, we additionally evaluate on scenes from BlendSwap [azinovic2022neural] and Scenethesis [ling2025scenethesis], which yield 26 test scenes including small on large relations and outdoor settings. iii) Real-world /stylized. For in-the-wild evaluation without ground truth, we test on images from DL3DV-140 [ling2024dl3dv], Gen3DSR [ardelean2024gen3dsr], ScanNet++ [yeshwanth2023scannet++], and stylized scenes. Then, we report qualitative comparisons.
@@ 4.2 Quantitative Results
Table 1 reports quantitative comparisons on synthetic datasets, including 3D-FRONT [fu20213d], BlendSwap [azinovic2022neural], and Scenethesis [ling2025scenethesis]. 3D-FRONT dataset only contains bedroom/living room scenarios, while Blendswap and Scenethesis contain more diverse layout. Our method, I-Scene, achieves the best performance among the state-of-the-art methods across all evaluated metrics without incurring much time consumption.
Geometry quality. Object-level. On 3D‑FRONT test scenes (ID), our method attains the lowest object‑level CD and the highest F‑score compared with MIDI [huang2025midi], SceneGen [meng2025scenegen], PartCrafter [lin2025partcrafter], and Gen3DSR [ardelean2024gen3dsr], indicating higher‑fidelity per‑object geometry. Crucially, on out-of-domain (OOD) benchmarks (BlendSwap, Scenethesis), our object‑level scores remain comparable to ID and still exceed all baselines by a clear margin, evidencing robust generalization. Scene-level. Across both ID and OOD settings, our method also achieves lowest scene‑level CD and highest F‑score than the baselines, reflecting more accurate and more complete whole‑scene geometry. We attribute these gains to: i) reprogrammed a pre-trained instance prior. A pre-trained instance generator supplies strong single‑object shape priors, which is an advantage over methods rely primarily on reconstruction from limited data such as Gen3DSR; ii) model‑centric supervision with shared, view‑centric scene context. Per‑instance generation conditioned on shared, view‑centric scene tokens keeps shapes near canonical while aligning pose and contact, reducing fused‑mesh artifacts and “shape bending” common in scene dataset‑supervised methods such as MIDI, SceneGen, PartCrafter. Unlike all baselines, which degrade substantially on OOD scenes, our object‑ and scene‑level metrics on BlendSwap and Scenethesis remain close to our ID scene performance, underscoring strong generalization.
@@ 4.4 Non-semantic random scene experiment
These findings indicate that i) I-Scene is able to learn spatial knowledge from non‑semantic scenes. In particular, instead of the high-level scene semantics, geometric cue alone provides a strong training signal for spatial learning and reasoning. ii) Curated scene annotations remain useful for in‑distribution calibration, synergizing training with non-semantic scenes to yield the best overall performance. and iii) Non‑semantic spatial relations are a valuable supervision source; scaling non‑semantic data provides richer spatial coverage that the model can exploit.
@@ 8 Experiment
Previous work MIDI [huang2025midi] uses ICP for alignment before calculating the metrics. We notice this ICP is not robust and leads to many bad alignment, making the metric calculation not reliable. Instead, we propose a robust ICP to make the metric results more reliable.
@@
Figure 6: We present the multi-view evaluation results using real-world /stylized image as input (example 1). The testing scenes contain various layouts including front, back, right, left, small on large, behind, and etc. Baselines includes PartCfrater [lin2025partcrafter], Gen3DSR [ardelean2024gen3dsr], MIDI-3D [huang2025midi], and SceneGen [meng2025scenegen]
Figure 7: We present the multi-view evaluation results using real-world /stylized image as input (example 2). The testing scenes contain various layouts including front, back, right, left, small on large, behind, and etc. Baselines includes PartCfrater [lin2025partcrafter], Gen3DSR [ardelean2024gen3dsr], MIDI-3D [huang2025midi], and SceneGen [meng2025scenegen]
@@
Figure 8: We present the multi-view evaluation results using synthetic scenes as input. The testing scenes contain various layouts including front, back, right, left, small on large, behind, and etc. Baselines includes PartCfrater [lin2025partcrafter], Gen3DSR [ardelean2024gen3dsr], MIDI-3D [huang2025midi], and SceneGen [meng2025scenegen]

### v2

@@ 1 Introduction
Generalization remains the central obstacle to interactive 3D scene generation. An effective system must produce editable, affordance‑aware, and spatially coherent object arrangements, capabilities critical for virtual content creation, simulation, and embodied AI. Recent end‑to‑end approaches extend powerful image‑to‑3D instance priors [24, 22] to multiple objects and relations in a single pass, yet their spatial understanding is typically learned from curated scene datasets whose limited coverage does not scale and ultimately constrains generalization to new layouts.
Recent methods learn layouts directly from annotated scenes [26, 21, 18, 8]; for example, SceneGen [18] explicitly models poses on 3D‑FRONT [5], and MIDI [8] learns inter‑object relations from those annotations. However, available interactive scene datasets are limited in scale, diversity, and spatial variety. For instance, the widely used interactive 3D scene dataset 3D-FRONT contains only $\sim$20K indoor bedroom and living-room scenes and under-represents small and supporting objects. Consequently, scene layout‑supervised models often overfit to dataset‑specific biases and fail to generalize to broader scene distributions—for example, small objects placed on or behind large furniture or outdoor settings.
@@ 2 Related Work
Single geometry 3D scene generation. One line of work represents entire scenes as a single global structure guided by text or visual prior. For example, Wonderland [13] predicts 3D Gaussian Splatting in a feed-forward manner from a single image; WonderWorld [30] generates connected scenes from a single view using layered Gaussian surfels. Similarly, WorldExporer [20]
@@
Compositional 3D interactive scene generation. Recent pipelines decompose scenes into perception and assembly: instance detection/segmentation and depth provide amodal cues, then objects are retrieved or generated, finally are assembled into a scene. Gen3DSR [1] follows a modular divide-and-conquer strategy; Deep Prior Assembly [31] integrates frozen priors (e.g., SAM [12], diffusion, Shap-E [11]) with layout fitting; REPARO [6] performs compositional asset generation with differentiable 3D layout alignment; CAST [28] applies component-aware analysis with SDF-based physics correction. On the text-driven side, graph/LLM planners (e.g., LayoutGPT [4], Holodeck [27], GALA3D [32]) synthesize scene graphs or relational constraints that downstream stages realize with assets and solvers. Scenethesis [15] and SceneWeaver [25] combines LLM planning, vision guidance, and physics-aware optimization with a plausibility judge. These approaches provide high fidelity and open-ended semantics, but are sensitive to early perception/planning errors, potential retrieval/solver mismatch, and incur per-scene optimization, limiting throughput and scalability.
Learning-based 3D interactive scene generation. Most recent approaches learn spatial layout distributions from curated 3D scene datasets [5], differing mainly in how they instantiate object assets. i) Layout-first scene synthesis. model a room as an unordered set of object tokens with attributes such as category, size, pose, which are later instantiated as geometry via retrieval or generation. For instance, ATISS [19], SceneFormer [23]) DiffuScene [21], MiDiffusion [7], DeBaRA [17], and PhyScene [26] condition on room type or floor plan, enabling controllable scene completion and synthesis; ii) End-to-end multi-instance method. With the advancement of pre-trained image-to-3D object model, recent approaches jointly model multiple objects and their spatial relations, reducing cascade errors and avoiding retrieval/solver loops at test time. MIDI-3D  [8] and SceneGen [18] model multiple assets and their implicit inter-object relations or instance explicit poses in one feed-forward pass. PartCrafter [14] extends compositional latent diffusion transformers to jointly denoise parts/objects into explicit triangle meshes. Both paradigms achieve strong control and coherence but remain constrained by the biases and limited diversity of existing scene datasets, often struggling with rare layouts or complex arrangements such as small objects supported by or occluded behind larger ones.
@@ 3.2 Overview
Figure 2 presents the overview of our generalizable spatial learner(I-Scene). I-Scene reprograms an image-to-3D instance foundation model to a spatial leaner. We use TRELLIS [24] backbone in the experiment. I-Scene only modifies the sparse structure transformer. Other stages are kept the same. I-Scene has two branches: spatial guidance branch and instance generation branch. The two branches share the same weight and are trained jointly.
@@ 3.4 View-Centric(VC) Space
Existing relevant works [8, 18] follow image-to-3D base model and use canonical space for scene representation. We observe using canonical space limits the learning of $F$ and hurts the model generalization ability. As illustrated in Figure 3 (a), given two camera ($C1$ and $C2$) and the scene composed of a blue square and a red circle in the space, the camera views of the scene are different as they are in different location and orientation, but the instance in the scene stay at the same position in the canonical space. So this formulation makes the $F$ ignores the object spatial position in the view image $I$ and only focuses on the local shape of each instance in the image. When training on some indoor dataset like 3D-FRONT, it is not significantly hurting $F$ as there are limited object in the scene and each instance usually has very different shapes. But if the scene has identical objects, e.g. several identical chairs having similar poses, $F$ often put duplicated chairs at the same position. We argue that the object spatial layout in the RGB image is also important and provides strong hint on its 3D layout pose. So we propose view-centric space where the space axis is based on the camera pose. As shown in Figure 3 (a), given the same two camera setting, the spatial layout of the two objects follows the change of the camera pose coherently. When training I-Scene in VC space, $F$ achieves much better generalization results as shown in our ablation study(Table 3).
@@ 3.5 Non-Semantic Synthetic 3D Scene
Collision-free Random Layout. We generate our synthetic scenes via sampling high‑quality 3D instances from a diverse 3D asset dataset(e.g., Objaverse [3]), randomly places them with a collision‑free mechanism to reduce severe occlusions, and enforces a common spatial relation such as right, left, front, back, and on the top of. The scene layout is purely non‑semantic meaningful composition of instance; only geometric and basic physical plausibility constraints are applied. By training on this dataset, I-Scene learns general spatial reasoning while remaining agnostic to category semantics. We provide our random layout generation details in appendix.
@@ 4.1 Experiment Setting
Baseline. We compare the representative interactive 3D scene generation methods including MIDI [8], SceneGen [18], Partcrafter [14], and Gen3DSR [1].
For all baselines, we use the same scene RGB image and instance masks as inputs, except PartCrafter, which does not support mask control; for PartCrafter we provide the scene RGB plus the target instance count. Because PartCrafter lacks texture rendering, we evaluate it only on scene-level geometry and layout quality. For visual quality comparisons, we texture MIDI’s generated scenes using MV-Adapter [9].
@@
Evaluation Dataset. We evaluate all methods on synthetic data and real-world/stylized scenes. i) Synthetic (in-domain). Following MIDI, we adopt the 3D-FRONT split and use the same test scenes, filtering out renders with fully occluded instance masks; this yields $\sim$860 test scenes. ii) Synthetic (out-of-domain). To assess generalization to novel layouts and object sets, we additionally evaluate on scenes from BlendSwap [2] and Scenethesis [15], which yield 26 test scenes including small on large relations and outdoor settings. iii) Real-world /stylized. For in-the-wild evaluation without ground truth, we test on images from DL3DV-140 [16], Gen3DSR [1], ScanNet++ [29], and stylized scenes. Then, we report qualitative comparisons.
@@ 4.2 Quantitative Results
Table 1 reports quantitative comparisons on synthetic datasets, including 3D-FRONT [5], BlendSwap [2], and Scenethesis [15]. 3D-FRONT dataset only contains bedroom/living room scenarios, while Blendswap and Scenethesis contain more diverse layout. Our method, I-Scene, achieves the best performance among the state-of-the-art methods across all evaluated metrics without incurring much time consumption.
Geometry quality. Object-level. On 3D‑FRONT test scenes (ID), our method attains the lowest object‑level CD and the highest F‑score compared with MIDI [8], SceneGen [18], PartCrafter [14], and Gen3DSR [1], indicating higher‑fidelity per‑object geometry. Crucially, on out-of-domain (OOD) benchmarks (BlendSwap, Scenethesis), our object‑level scores remain comparable to ID and still exceed all baselines by a clear margin, evidencing robust generalization. Scene-level. Across both ID and OOD settings, our method also achieves lowest scene‑level CD and highest F‑score than the baselines, reflecting more accurate and more complete whole‑scene geometry. We attribute these gains to: i) reprogrammed a pre-trained instance prior. A pre-trained instance generator supplies strong single‑object shape priors, which is an advantage over methods rely primarily on reconstruction from limited data such as Gen3DSR; ii) model‑centric supervision with shared, view‑centric scene context. Per‑instance generation conditioned on shared, view‑centric scene tokens keeps shapes near canonical while aligning pose and contact, reducing fused‑mesh artifacts and “shape bending” common in scene dataset‑supervised methods such as MIDI, SceneGen, PartCrafter. Unlike all baselines, which degrade substantially on OOD scenes, our object‑ and scene‑level metrics on BlendSwap and Scenethesis remain close to our ID scene performance, underscoring strong generalization.
@@ 4.4 Non-semantic random scene experiment
These findings indicate that i) I-Scene is able to learn spatial knowledge from non‑semantic scenes. In particular, rather than relying on high-level scene semantics, pure geometric cues alone provide a strong supervision signal for spatial learning and reasoning. This observation suggests a broader scaling law: analogous to MegaSynth [10] in 3D reconstruction, scaling synthetic, non-semantic layouts would provide richer spatial coverage that the model can exploit thus unlock large-scale interactive 3D scene generation. ii) Curated scene annotations remain useful for in‑distribution calibration, synergizing training with non-semantic scenes to yield the best overall performance.
@@ 8 Experiment
Previous work MIDI [8] uses ICP for alignment before calculating the metrics. We notice this ICP is not robust and leads to many bad alignment, making the metric calculation not reliable. Instead, we propose a robust ICP to make the metric results more reliable.
@@
Figure 6: We present the multi-view evaluation results using real-world /stylized image as input (example 1). The testing scenes contain various layouts including front, back, right, left, small on large, behind, and etc. Baselines includes PartCfrater [14], Gen3DSR [1], MIDI-3D [8], and SceneGen [18]
Figure 7: We present the multi-view evaluation results using real-world /stylized image as input (example 2). The testing scenes contain various layouts including front, back, right, left, small on large, behind, and etc. Baselines includes PartCfrater [14], Gen3DSR [1], MIDI-3D [8], and SceneGen [18]
@@
Figure 8: We present the multi-view evaluation results using synthetic scenes as input. The testing scenes contain various layouts including front, back, right, left, small on large, behind, and etc. Baselines includes PartCfrater [14], Gen3DSR [1], MIDI-3D [8], and SceneGen [18]